OpenStudy (happycakes):

Solve the following system of equations enter the y coordinate of the solution.5x+2y=7 -2x+6y=9

OpenStudy (will.h):

Hey Welsh. Hint solve either by substitution method or elimination

OpenStudy (welshfella):

use elimination 5x + 2y = 7 multiply this by -3 -15x - 6y = -21 ----- (1) -2x + 6y = 9 ------- (2) adding -17x = -12 now find x then you can find y by plugging in this value for x into the first equation and solving for y

OpenStudy (will.h):

Legend bruh

OpenStudy (welshfella):

Hi Will

OpenStudy (will.h):

Doing a great job as always :)

OpenStudy (welshfella):

ty So you are 18 y o Wish I was lol!

OpenStudy (will.h):

Lol are you that old!?